A Level Mathematics 9709 Paper 52, Oct/Nov 2024

Question paper and mark scheme50 marks7 questions

Solve it on ExamPal

Practise

Untimed. Write on the paper and check as you go.

Start

Exam

Timed, with a clock you set, like sitting the real paper.

Start

Opens full screen.

Papers open in

Opening the question paper

Read as textThe questions and the mark scheme as plain text. Easier to read on a phone.

Question paper

9709/52 · Oct/Nov 2024

Download PDF
1

At a college, the students choose exactly one of tennis, hockey or netball to play. The table shows the numbers of students in Year 1 and Year 2 at the college playing each of these sports.

TennisHockeyNetball
Year 1162212
Year 2241828

One student is chosen at random from the 120 students. Events X and N are defined as follows: X: the student is in Year 1 N: the student plays netball.

a

Find P(X∣N).

[1]
b

Find P(N∣X).

[1]
c

Determine whether or not X and N are independent events.

[1]
d

One of the students who plays netball takes 8 shots at goal. On each shot, the probability that she will succeed is 0.15, independently of all other shots. Find the probability that she succeeds on fewer than 3 of these shots.

[3]
2
a

Find the number of different arrangements of the 9 letters in the word ALGEBRAIC.

[1]
b

Find the number of different arrangements of the 9 letters in the word ALGEBRAIC in which there are no more than two letters between the two As.

[3]
3

A fair coin and an ordinary fair six-sided dice are thrown at the same time. The random variable X is defined as follows. • If the coin shows a tail, X is twice the score on the dice. • If the coin shows a head, X is the score on the dice if the score is even and X is 0 otherwise.

a

Draw up the probability distribution table for X.

[3]
b

Find Var⁡(X).

[3]
4

The heights, in metres, of white pine trees are normally distributed with mean 19.8 and standard deviation 2.4 . In a certain forest there are 450 white pine trees.

a

How many of these trees would you expect to have height less than 18.2 metres?

[4]
b

The heights, in metres, of red pine trees are normally distributed with mean 23.4 and standard deviation σ. It is known that 26% of red pine trees have height greater than 25.5 metres. Find the value of σ.

[3]
5

In a class of 21 students, there are 10 violinists, 6 guitarists and 5 pianists. A group of 7 is to be chosen from these 21 students. The group will consist of 4 violinists, 2 guitarists and 1 pianist.

a

In how many ways can the group of 7 be chosen?

[2]
b

On another occasion a group of 5 will be chosen from the 21 students. The group must contain at least 2 violinists, at least 1 guitarist and at most 1 pianist. In how many ways can the group of 5 be chosen?

[4]
6

Teams of 15 runners took part in a charity run last Saturday. The times taken, in minutes, to complete the course by the runners from the Falcons and the runners from the Kites are shown in the table.

Falcons383942444648505152565859646976
Kites324040454748525458595960616365
a

Draw a back-to-back stem-and-leaf diagram to represent this information, with the Falcons on the left-hand side.

[4]
b

Find the median and the interquartile range of the times for the Falcons.

[3]
c

Let x and y denote the times, in minutes, of a runner from the Falcons and a runner from the Kites respectively. It is given that ∑x=792,∑x2=43 504,∑y=783,∑y2=42 223. Find the mean and the standard deviation of the times taken by all 30 runners from the two teams.

[3]
7

In a game, players attempt to score a goal by kicking a ball into a net. The probability that Leno scores a goal is 0.4 on any attempt, independently of all other attempts. The random variable X denotes the number of attempts that it takes Leno to score a goal.

a

Find P(X=5).

[1]
b

Find P(3⩽X⩽7).

[2]
c

Find the probability that Leno scores his second goal on or before his 5th attempt.

[3]
d

Leno has 75 attempts to score a goal. Use a suitable approximation to find the probability that Leno scores more than 28 goals but fewer than 35 goals.

[5]

Mark scheme

9709/52 · Oct/Nov 2024

Download PDF

1a · 1 mark

  • B1

    [P(X|N) =] 12/40

1b · 1 mark

  • B1

    [P(N|X) =] 12/50

1c · 1 mark

  • B1

    P(N ∩ X) = 12/120, P(N) = 40/120, P(X) = 50/120 40/120 × 50/120 = 5/36, 0.138[8...] ≠ 12/120, 0.1 Not independent

1d · 3 marks

  • M1

    One term of form ^8C_x (p)^x (1 – p)^(8–x) with 0 < p < 1, x ≠ 0 or 8

  • A1

    Correct unsimplified expression, no terms omitted leading to final answer

  • B1

    = 0.895

2a · 1 mark

  • B1

    [9! / 2! =] 181440

2b · 3 marks

  • B1

    Correct outcome/value for 1 identified scenario, accept unsimplified

  • M1

    Add values of 3 correct scenarios, no incorrect/repeated scenarios (or their 2(a), or correct, subtract values of 5 correct scenarios, no incorrect/repeated scenarios)

  • A1

    = 105 840

3a · 3 marks

  • B1

    Table with correct x values and at least 2 correct probabilities

  • B1

    Four more probabilities correctly linked to the correct x value, need not be in table, accept unsimplified

  • B1

    7 correct probabilities linked with correct outcomes, may not be in table

3b · 3 marks

  • M1

    [E(X) =] [0 +] 2 × 2/12 + 4 × 2/12 + 6 × 2/12 + 8 × 1/12 + 10 × 1/12 + 12 × 1/12

  • M1

    Appropriate variance formula using their (E(X))^2 value

  • A1

    = 14.75, 14 3/4

4a · 4 marks

  • M1

    Use of ± standardisation formula with 18.2, 19.8 and 2.4 substituted appropriately, no continuity correction

  • M1

    Calculating the appropriate probability areas (leading to their final answer, expect < 0.5)

  • A1

    = 0.252(3)

  • 1 mark

    [Expected number = 0.2523 × 450 = 113.5, ] = 113 or 114

4b · 3 marks

  • B1

    0.643 ⩽ z ⩽ 0.6435 or −0.6435 ⩽ z ⩽ −0.643 seen

  • M1

    ± standardisation formula with 25.5, 23.4, σ equating to a z-value, (not 1 – their z-value…)

  • A1

    σ = 3.27

5a · 2 marks

  • M1

    ^10C_a × ^6C_b × ^5C_c , a + b + c = 7, a,b,c integers

  • A1

    [= 210 × 15 × 5] = 15750

5b · 4 marks

  • M1

    One product using 2 or 3 combinations with upper numbers correct and lower numbers summing to 5 and linked to a correct identified scenario

  • B1

    2 identified outcomes evaluated accurately, accept unsimplified

  • M1

    Add values of 5 correct scenarios, no incorrect/repeated scenarios

  • A1

    Total = 10935

6a · 4 marks

  • B1

    Correct stem, ignore extra values (not in reverse, not split)

  • B1

    Correct Falcons labelled on left, leaves in order from right to left and lined up vertically, no commas or other punctuation

  • B1

    Correct Kites labelled on same diagram, leaves in order and lined up vertically, no commas or other punctuation

  • B1

    Correct key, for their diagram, need both teams names and ‘mins’ at least once here, or in leaf headings or title

6b · 3 marks

  • B1

    Median = 51 [minutes]

  • M1

    [IQR =] 59 – 44

  • A1

    = 15 [minutes]

6c · 3 marks

  • B1

    [mean = (792 + 783)/30 = 1575/30 =] 52.5, 52 1/2, 105/2

  • M1

    Accept unsimplified variance formula. FT their mean. Ignore any square root leading to sd for this mark

  • A1

    σ = (√101.3167 =) 10.1

7a · 1 mark

  • B1

    [(0.6)^4 × 0.4 =] 0.0518[4], 162/3125

7b · 2 marks

  • M1

    (1 – p^7) – (1 – p^2) or p^2 – p^7 seen, 0 < p < 1 (or (1 – p) × p^2 + (1 – p) × p^3 + (1 – p) × p^4 + (1 – p) × p^5 + (1 – p) × p^6 seen, 0 < p < 1 or 0.144(1 – p^5)/(1 – p) seen 0 < p < 1)

  • A1

    = 0.332[0... ], 25938/78125

7c · 3 marks

  • M1

    2 correct unsimplified outcomes (or at least 2 correct unsimplified terms)

  • M1

    Add values for 4 identified correct scenarios (or add values for 4 terms of the form ^5C_a(0.4)^a(0.6)^(5–a) or 1 – sum of 2 terms of the form ^5C_a(0.4)^a(0.6)^(5–a))

  • A1

    = 0.663, 2072/3125

7d · 5 marks

  • B1

    [Mean = 75 × 0.4 =] 30 [Variance = 75 × 0.4 × 0.6 =] 18

  • M1

    Substituting their µ and positive σ into one ±standardising formula (any number for 28.5 or 34.5), not σ^2, not √σ

  • M1

    Using continuity corrections 27.5 or 28.5 and 34.5 or 35.5 in their 2 separate standardisation formula

  • M1

    Appropriate area Φ, from final process. Must be a probability

  • A1

    = 0.494

Questions

  1. 1
    Question 16 marks
    Mark scheme ›
  2. 2
    Question 24 marks
    Mark scheme ›
  3. 3
    Question 36 marks
    Mark scheme ›
  4. 4
    Question 47 marks
    Mark scheme ›
  5. 5
    Question 56 marks
    Mark scheme ›
  6. 6
    Question 610 marks
    Mark scheme ›
  7. 7
    Question 711 marks
    Mark scheme ›

More 9709 papers

Other papers from Oct/Nov 2024

Solve it on ExamPalStart